OpenStudy (anonymous):

For a closed cylinder with radius r cm and height h cm, find the dimensions giving the minimum surface area, given that the volume is 8 cm3. Write the EXACT answers. r=_____ h=_____ please explain? :) thanks!! @Luigi0210

OpenStudy (anonymous):

I use these right? V= π * r^2 * h and A= π * r * (2h+r) ?

OpenStudy (anonymous):

and do i set it like this? 8=π * r^2 * h h=8/π * r^2 ?

OpenStudy (agent0smith):

Differentiate the SA equation, set it equal to 0 to find where it's a minimum.

OpenStudy (anonymous):

it's a closed cylinder, not an opened one. Your surface area formula is incorrect

OpenStudy (anonymous):

ahh okay darn.. so how would we solve this problem then?

OpenStudy (agent0smith):

Yeah it should be \[\large SA = 2 \pi r h + 2 \pi r^2\]

OpenStudy (anonymous):

A = 2pi r^2 + 2pi r h = 2pi r (r + h)

OpenStudy (anonymous):

ohhh okay so then you set it equal to 8?

OpenStudy (anonymous):

and V = pi r^2 h

OpenStudy (anonymous):

so is it 8=pi r^2 h so solving for h, h= 8 / pi r^2 ?

OpenStudy (anonymous):

yes V = pi r^2 h

OpenStudy (anonymous):

8 = pi r^2 h

OpenStudy (anonymous):

so solving for h right? h= 8 -------- π r^2 ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

and then set A' = 0, and solve for r. Once you have r, you can find h.

OpenStudy (anonymous):

A' ? the derivative of 2pi r (r + h) ?

OpenStudy (anonymous):

not sure what to do here :/

OpenStudy (anonymous):

@sourwing ? @agent0smith ?

OpenStudy (agent0smith):

Plug that equation for h into the SA formula. Then differentiate w.r.t. r.

OpenStudy (anonymous):

wrtr?

OpenStudy (anonymous):

so 2πrh+2πr^2 = 2πr(8/π r^2) + 2 π r^2 ?

OpenStudy (anonymous):

= 16πr + 2π r^2 ----------------- π r^2 ?

OpenStudy (anonymous):

correct. Then find dA/dr

OpenStudy (anonymous):

the derivative?

OpenStudy (anonymous):

how can I do that? i'm not quite sure :(

OpenStudy (anonymous):

@sourwing ?

OpenStudy (anonymous):

A = 2πr(8/π r^2) + 2 π r^2 A = 16 / r + 2pi r^2 A' = -16/r^2 + 4pi r

OpenStudy (anonymous):

ohhh okay i see so what do we do from here?

OpenStudy (anonymous):

set it equal to 0 and solve for r

OpenStudy (anonymous):

okay so -16 ---- + 4πr = 0 r^2 -16/r^2 = -4πr -16 = -4πr * r^2 -16 = -4πr^3 4/π = r^3 ? did I do that right so far? @sourwing :/

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so r=4/π^(1/3)

OpenStudy (anonymous):

no, r = (4/pi)^(1/3)

OpenStudy (anonymous):

ahh okay:) so that's our r value right? so now we need to find the h?

OpenStudy (anonymous):

you had the formula for h, find it, plug r in

OpenStudy (anonymous):

so 8 ----- pi ((4/pi)^(1/3))^2 ? not sure how to simplify that part though :/

OpenStudy (anonymous):

@sourwing ?

OpenStudy (anonymous):

8 -- pi (4/pi)^(2/3)

OpenStudy (anonymous):

oh! and that's as simplified as it gets?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

ahh okay awesome!!Thank you so much!! :D